7. An anti aircraft missile is fired vertically upward from the earth’s surface with an initial

velocity of 500 m/s. Calculate:

(a) The maximum height it can reach.

(b) The time taken to reach that height.

(c) The velocity of the missile at t = 60 s.

(d) The time when it reaches the height of 10 km.


1
Expert's answer
2021-11-16T09:58:46-0500

(a) The maximum height it can reach:


"H_\text{max}=\frac{v^2}{2g}=12755\text{ m}."

(b) The time taken to reach that height:


"t=\frac{v}{g}=51\text{ s}."

(c) The velocity of the missile at t = 60 s:

60 s is 9 seconds from the max. height, so, the velocity is


"v_c=gt=9.8·9=88.2\text{ m/s}."

(d) The time when it reaches the height of 10 km.


"\tau=t-\sqrt{2(y)/g},\\
\tau=51-\sqrt{2(2755)/9.8}=27.3\text{ s}."